# Convergence of empirical Gromov-Wasserstein distance

### v1

@@ 3.2. Minimax lower bounds
The proof builds on adapting techniques developed in [NWR22] (see also [MNW24]) and compares $\mathsf{GW}_{2,2}$ with the total vatiation and the product of the total variaton and $\chi^{2}$-divergence for random mixtures of point masses; see (21). The upper bound is straightforward from comparing $\mathsf{GW}_{2,2}$ with $W_{2}$ using Lemma 2.1. Deriving the lower comparison inequality requires more effort. First, to apply the second result in Lemma 2.1 (ii), one has to appropriately choose support points so that they are well-separated and at the same time the smallest eigenvalue of the covariance matrix of the empirical distribution is bounded away from zero; see Lemma 4.4. Second, extra care is needed to bound $\mathsf{GW}_{2,2}$ from below by the total variation because of invariance of $\mathsf{GW}_{2,2}$ under isometries; see the argument above (21).
@@ 4. Proofs
We will use the following observation without further mentioning: $m_{p}(\mu)\leq 1+m_{p^{\prime}}(\mu)$ for any $p<p^{\prime}$.
@@
$\sup_{\begin{subarray}{c}(\mu,\nu):~m_{4q}(\mu)\vee m_{4q}(\nu)\leq M\\ \|A\|_{\operatorname{op}}\leq R\end{subarray}}\mathbb{E}\left[\big{|}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{|}\right]\lesssim\varphi_{n,m},$
where the hidden constant depends only on $q,d_{x},d_{y},R$ and $M$.
@@
The proof is essentially contained in the proof of Theorem 4.2 in [ZGMS24], so we only provide an outline. For any random vector $(X,Y)$ supported in $B_{\mathcal{X}}(0,r)\times B_{\mathcal{Y}}(0,r)$,
@@
From $\|\cdot\|_{\infty}$-entropy number estimates for $\mathcal{F}$ (cf. Corollary 2.7.10 in [vdVW96]), combined with a version of Dudley’s entropy integral bound (cf. Lemma 16 in [vLB04]), we have
$\mathbb{E}[I]\lesssim\inf_{\alpha>0}\left[\alpha+\frac{1}{\sqrt{n}}\int_{\alpha}^{1}\tau^{-d_{x}/4}\,d\tau\right]\lesssim n^{-\frac{2}{d_{x}\vee 4}}(\log n)^{\mathbbm{1}_{\{d_{x}=4\}}}$
up to a constant that depends only on $d_{x},d_{y}$ and $R$. For the second term $II$, by Lemma 2.1 in [HSM24] (or by the definition of the $c_{A}$-transform), $\|\cdot\|_{\infty}$-covering numbers for $\mathcal{F}^{c_{A}}$ are no greater than those of $\mathcal{F}$, and as such, arguing as in the previous case, we have $\mathbb{E}[II]\lesssim m^{-\frac{2}{d_{x}\vee 4}}(\log m)^{\mathbbm{1}_{\{d_{x}=4\}}}$ up to a constant that depends only on $d_{x},d_{y}$ and $R$. ∎
@@
$\begin{split}&\mathbb{E}\left[\big{|}T_{\bar{c}_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{\bar{c}_{A}}(\mu,\nu)-T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})+T_{c_{A}}(\mu,\nu)\big{|}\right]\\ &\leq\mathbb{E}\Big{[}4\big{|}m_{4}(\hat{\mu}_{n})-m_{4}(\mu)\big{|}+4\big{|}m_{4}(\hat{\nu}_{m})-m_{4}(\nu)\big{|}\\ &\qquad+16R\big{|}\mu_{2}(\hat{\mu}_{n})-m_{2}(\mu)\big{|}+16R\big{|}m_{2}(\hat{\nu}_{m})-m_{2}(\nu)\big{|}\Big{]}\\ &\leq\left\{4\sqrt{m_{8}(\mu)}+4\sqrt{m_{8}(\nu)}+16R\sqrt{m_{4}(\mu)}+16R\sqrt{m_{4}(\nu)}\right\}(n\wedge m)^{-1/2}.\end{split}$
@@
$\begin{split}\bar{c}_{A}(x,y)&\leq(8\|x\|^{4}+32R\|x\|^{2})+(8\|y\|^{4}+32R\|y\|^{2})\\ &\leq\underbrace{\big{\{}(8+32R)\|x\|^{4}+32R\big{\}}}_{=:\bar{c}_{\mathcal{X}}(x)}+\underbrace{\big{\{}(8+32R)\|y\|^{4}+32R\big{\}}}_{=:\bar{c}_{\mathcal{Y}}(y)},\end{split}$
and $\|\bar{c}_{\mathcal{X}}\|_{L^{q}(\mu)}^{q}\vee\|\bar{c}_{\mathcal{Y}}\|_{L^{q}(\nu)}^{q}\lesssim 1$ up to a constant that depends only on $q,R$ and $M$. Since
$\bar{c}_{\mathcal{X}}^{-1}([0,r])=\begin{cases}B_{\mathcal{X}}\left(0,\left(\frac{r-32R}{8+32R}\right)^{1/4}\right)&r\geq 32R,\\ \varnothing&r<32R,\end{cases}$
Condition (22) holds with
$\alpha=\frac{2}{(d_{x}\wedge d_{y})\vee 4}\quad\text{and}\quad\delta=1.$
The desired result then follows from Theorem A.1. ∎
@@
Let $q,R\geq 1$ be given. Set $\bar{c}_{\mathcal{X}}(x):=(8+32R)\|x\|^{4}+32R$ and $\bar{c}_{\mathcal{Y}}(y):=(8+32R)\|y\|^{4}+32R$. For every $(\mu,\nu)\in\mathcal{P}_{4q}(\mathcal{X})\times\mathcal{P}_{4q}(\mathcal{Y})$ and $A\in\mathbb{R}^{d_{x}\times d_{y}}$ with $\|A\|_{\operatorname{op}}\leq R$, one can find dual potentials $(f_{A},g_{A})$ for $(\mu,\nu)$ w.r.t. cost $c_{A}$ such that
@@
$f_{A}(x):=\bar{f}_{A}(x)+(4\|x\|^{4}+16R\|x\|^{2})\quad\text{and}\quad g_{A}(y):=\bar{g}_{A}(y)+(4\|y\|^{4}+16R\|y\|^{2})$
@@
Pick any $(\mu,\nu)\in\mathcal{P}_{4q}(\mathcal{X})\times\mathcal{P}_{4q}(\mathcal{Y})$ with $m_{4q}(\mu)\vee m_{4q}(\nu)\leq M$. In this proof, the notation $\lesssim$ means that an inequality holds up to a constant that depends only on $q,d_{x},d_{y}$ and $M$. Assume without loss of generality that $\mu$ and $\nu$ have mean zero. Let $\tilde{\mu}_{n}$ and $\tilde{\nu}_{m}$ be the centered versions of $\hat{\mu}_{n},\hat{\nu}_{m}$, respectively. We first observe that
@@
These estimates follow from tedious but straightforward computations, having in mind, e.g., $\mathbb{E}[\|\bar{X}\|^{4}]\lesssim n^{-2}$ for $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$.
Let $A^{\star}$ be a matrix with $\|A^{\star}\|_{\operatorname{op}}\leq\sqrt{m_{2}(\mu)m_{2}(\nu)}/2$ that achieves the infimum in the variational representation (6),
@@
$\mathcal{E}_{n,m}=\big{\{}m_{2}(\hat{\mu}_{n})\leq m_{2}(\mu)+1\big{\}}\cap\{m_{2}(\hat{\nu}_{m})\leq m_{2}(\nu)+1\big{\}}.$
By Markov’s inequality, we have
$\begin{split}\mathbb{P}(\mathcal{E}_{n,m}^{c})&\leq\mathbb{P}\big{(}m_{2}(\hat{\mu}_{n})-m_{2}(\mu)\geq 1\big{)}+\mathbb{P}\big{(}m_{2}(\hat{\nu}_{m})-m_{2}(\nu)\geq 1\big{)}\\ &\leq\frac{m_{4}(\mu)}{n}+\frac{m_{4}(\nu)}{m}\\ &\leq\big{(}m_{4}(\mu)+m_{4}(\nu)\big{)}(n\wedge m)^{-1}.\end{split}$
Observe that on the event $\mathcal{E}_{n}$, by Lemma 2.2, the following variational representation holds:
@@
where $\mathcal{A}:=\big{\{}A\in\mathbb{R}^{d_{x}\times d_{y}}:\|A\|_{\operatorname{op}}\leq R\big{\}}$ with $R:=\sqrt{(m_{2}(\mu)+1)(m_{2}(\nu)+1)}/2$. An analogous variational representation holds for $S_{2}(\mu,\nu)$. As such, on the event $\mathcal{E}_{n}$,
@@
$\begin{split}|c_{A}(x,y)-c_{B}(x,y)|&\leq 32|x^{\top}(A-B)y|\\ &\leq 32\|x\|\|y\|\|A-B\|_{\operatorname{op}}\\ &\leq 16(\|x\|^{2}+\|y\|^{2})\|A-B\|_{\operatorname{op}}.\end{split}$
For $\epsilon>0$, let $\mathcal{N}_{\epsilon}$ be an $\epsilon$-net for $\mathcal{A}$ w.r.t. $\|\cdot\|_{\operatorname{op}}$, so that for any $A\in\mathcal{A}$, there exists $B\in\mathcal{N}_{\epsilon}$ such that
$\|A-B\|_{\operatorname{op}}\leq\epsilon.$
By volumetric argument, it is not difficult to see that
@@
$\begin{split}\inf_{A\in\mathcal{A}}\big{\{}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{\}}&\geq\min_{A\in\mathcal{N}_{\epsilon_{n,m}}}\big{\{}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{\}}\\ &\quad-16\big{(}m_{2}(\hat{\mu}_{n})+m_{2}(\hat{\nu}_{m})+m_{2}(\mu)+m_{2}(\nu)\big{)}\epsilon_{n,m}.\end{split}$
@@
Outside $\mathcal{E}_{n}$, we use the following crude bound:
$\begin{split}S_{2}(\hat{\mu}_{n},\hat{\nu}_{m})-S_{2}(\mu,\nu)&\geq-4m_{4}(\hat{\mu}_{n})-16Rm_{2}(\hat{\mu}_{n})-4m_{4}(\hat{\nu}_{m})-16Rm_{2}(\hat{\nu}_{m})\\ &\quad-4m_{4}(\mu)-16Rm_{2}(\mu)-4m_{4}(\nu)-16Rm_{2}(\nu).\end{split}$
@@
$\begin{split}\mathbb{E}\left[\big{|}S_{2}(\hat{\mu}_{n},\hat{\nu}_{m})-S_{2}(\mu,\nu)\big{|}\right]&\lesssim\mathbb{E}\left[\big{|}T_{c_{A^{\star}}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A^{\star}}}(\mu,\nu)\big{|}\right]\\ &\qquad+\mathbb{E}\left[\max_{A\in\mathcal{N}_{\epsilon_{n,m}}}\left|\int f_{A}\,d(\hat{\mu}_{n}-\mu)\right|\right]\\ &\qquad+\mathbb{E}\left[\max_{A\in\mathcal{N}_{\epsilon_{n,m}}}\left|\int g_{A}\,d(\hat{\nu}_{m}-\nu)\right|\right]\\ &\qquad+\mathbb{E}\big{[}\big{(}m_{4}(\hat{\mu}_{n})+m_{4}(\hat{\nu}_{m})\big{)}\mathbbm{1}_{\mathcal{E}_{n,m}^{c}}\big{]}+(n\wedge m)^{-\frac{2}{(d_{x}\wedge d_{y})\vee 4}}.\end{split}$

### v2

@@ 3.2. Minimax lower bounds
The proof builds on adapting techniques developed in [NWR22] (see also [MNW24]) and compares $\mathsf{GW}_{2,2}$ with the total variation and the product of the total variation and $\chi^{2}$-divergence for random mixtures of point masses; see (29). The upper bound is straightforward from comparing $\mathsf{GW}_{2,2}$ with $W_{2}$ using Lemma 2.1. Deriving the lower comparison inequality requires more effort. First, to apply the second result in Lemma 2.1 (ii), one has to appropriately choose support points so that they are well-separated and at the same time the smallest eigenvalue of the covariance matrix of the empirical distribution is bounded away from zero; see Lemma 4.6. Second, extra care is needed to bound $\mathsf{GW}_{2,2}$ from below by the total variation because of invariance of $\mathsf{GW}_{2,2}$ under isometries; see the argument above (29).
Theorem 3.2 implies that the rate $\varphi_{n,m}$ is unimprovable (up to logarithmic factors) uniformly over the class of distributions $\mathcal{P}(B_{\mathcal{X}}(0,1))\times\mathcal{P}(B_{\mathcal{X}}(0,1))$ (or any larger distributional class). However, this does not preclude faster rates for specific distributions. For example, when $\mu,\nu$ are both the uniform distribution on $0,1]^{2}$, Lemma 2.1 implies that $D(\hat{\mu}_{n},\hat{\nu}_{n})=\mathsf{GW}_{2,2}^{2}(\hat{\mu}_{n},\hat{\nu}_{n})\lesssim\mathsf{GW}_{2,1}^{2}(\hat{\mu}_{n},\hat{\nu}_{n})\lesssim W_{2}^{2}(\hat{\mu}_{n},\hat{\nu}_{n})$. Theorem 1.1 in [AST19] shows
$\lim_{n\to\infty}\frac{n}{\log n}\mathbb{E}\big{[}W_{2}^{2}(\hat{\mu}_{n},\hat{\nu}_{n})\big{]}=\frac{1}{2\pi},$
which implies that $\mathbb{E}[D(\hat{\mu}_{n},\hat{\nu}_{n})]\lesssim\frac{\log n}{n}$. Exploring faster rates of GW for restricted classes of distributions is left for future research.
3.3. Deviation inequalities
In this section, we present deviation inequalities for the error of empirical GW. For the notation simplicity, we shall focus here on $m=n$. Define
$\Delta_{n}:=\big{|}D(\hat{\mu}_{n},\hat{\nu}_{n})-D(\mu,\nu)\big{|},\quad\varphi_{n}:=\varphi_{n,n}=n^{-\frac{2}{(d_{x}\wedge d_{y})\vee 4}}(\log n)^{\mathbbm{1}_{\{d_{x}\wedge d_{y}=4\}}},$
and $\bar{\varphi}_{n,q}$ by the right-hand side on (11) with $m=n$. Our goal is to establish deviation inequalities for $\Delta_{n}$. We consider the following three cases for the marginals $\mu,\nu$; (i) they are compactly supported, (ii) they are sub-Weibull, and (iii) they have finite $(4q)$-th moments for some $q\in(2,\infty)$. Recall that a real-valued random variable $\xi$ is called $\beta$-sub-Weibull for some $\beta>0$ if the Orlicz $\psi_{\beta}$-norm with $\psi_{\beta}(z):=e^{z^{\beta}}-1$ is finite, i.e.,
$\|\xi\|_{\psi_{\beta}}:=\inf\left\{C>0:\mathbb{E}\big{[}e^{|\xi/C|^{\beta}}\big{]}\leq 2\right\}<\infty.$
See [KC22] for sub-Weibull distributions. [^5] For $\beta\in(0,1)$, $\|\cdot\|_{\psi_{\beta}}$ is only a quasinorm. A $\beta$-sub-Weibull variable $\xi$ is often called sub-Gaussian if $\beta=2$ and sub-exponential if $\beta=1$. For a random vector $Z$, we call it $\beta$-sub-Weibull if $\|Z\|_{\psi_{\beta}}:=\|\|Z\|\|_{\psi_{\beta}}$ is finite.
Theorem 3.3 (Deviation inequalities for empirical GW error).
Let $r,\kappa,M\geq 1,\beta>0$ and $q\in(2,\infty)$ be given.
  1. (i)
    If $\mu,\nu$ are supported in $B_{\mathcal{X}}(0,r),B_{\mathcal{Y}}(0,r)$, respectively, then
    $\mathbb{P}\Big{(}\Delta_{n}\geq Kr^{4}\big{(}\varphi_{n}+tn^{-1/2}\big{)}\Big{)}\leq 2e^{-t^{2}},\quad\forall t>0,$
    where $K$ is a constant that depends only on $d_{x}$ and $d_{y}$.
  2. (ii)
    If $\mu,\nu$ are $\beta$-sub-Weibull with $\|X\|_{\psi_{\beta}}\vee\|Y\|_{\psi_{\beta}}\leq M$ for $X\sim\mu$ and $Y\sim\nu$, then
    $\begin{split}&\mathbb{P}\Big{(}\Delta_{n}\geq K\big{(}\varphi_{n}+n^{-1/2}\sqrt{\log n}+stn^{-1/2}+tn^{-1/2}(\log n)^{4/\beta}\big{)}\Big{)}\\ &\qquad\leq 2n^{-\kappa}e^{-s^{\beta/4}}+2e^{-t^{2}},\quad\forall s\geq 1,t>0,\end{split}$
    where $K$ is a constant that depends only on $d_{x},d_{y},\kappa,\beta$ and $M$.
  3. (iii)
    If $\mathfrak{m}_{4q}(\mu)\vee\mathfrak{m}_{4q}(\nu)\leq M$, then
    $\mathbb{P}\Big{(}\Delta_{n}\geq K\big{(}\bar{\varphi}_{n,q}+stn^{-\frac{1}{2}+\frac{1}{q}}+s^{-q+1}n^{1/q}\big{)}\Big{)}\leq 2s^{-q}+2e^{-t^{2}},\quad\forall s\geq 1,t>0,$
    where $K$ is a constant that depends only on $d_{x},d_{y},q$ and $M$.
The theorem implies that when $d_{x}\wedge d_{y}>4$, $\Delta_{n}$ can be bounded by a constant multiple of $\varphi_{n}$ with high probability, provided $q$ is large enough in Case (iii). Suppose $d_{x}\wedge d_{y}>4$.
  1. (i)
    When $\mu,\nu$ are compactly supported, $\Delta_{n}\lesssim n^{-2/(d_{x}\wedge d_{y})}$ with probability at least, say, $1-n^{-10}$. This follows by choosing $t=K^{\prime}\sqrt{\log n}$ for a sufficiently large $K^{\prime}$. [^6] We used $n\geq 2$ to eliminate the constant factor in front of $n^{-10}$.
  2. (ii)
    Likewise, when $\mu,\nu$ are sub-Weibull, $\Delta_{n}\lesssim n^{-2/(d_{x}\wedge d_{y})}$ with probability at least $1-n^{-10}$. This follows by choosing $s=K^{\prime}(\log n)^{4/\beta}$ and $t=K^{\prime}\sqrt{\log n}$ for a sufficiently large $K^{\prime}$.
  3. (iii)
    Suppose $\mu,\nu$ have finite $(4q)$-th moments for some $q$ large enough that $\bar{\varphi}_{n,q}\lesssim n^{-2/(d_{x}\wedge d_{y})}$. Choosing $s=n^{1/(2q)}$ and $t=K^{\prime}\sqrt{\log n}$ for a suitable constant $K^{\prime}$ yields $stn^{-\frac{1}{2}+\frac{1}{q}}+s^{-q+1}n^{1/q}\lesssim n^{-\frac{1}{2}+\frac{3}{2q}}\sqrt{\log n}\lesssim n^{-2/(d_{x}\wedge d_{y})}$, provided $q$ is large enough. As such, we have $\Delta_{n}\lesssim n^{-2/(d_{x}\wedge d_{y})}$ with probability at least $1-O(n^{-1/2})$.
In particular, in Cases (i) and (ii) above, by the Borel-Cantelli lemma, we have
$\limsup_{n\to\infty}n^{2/(d_{x}\wedge d_{y})}\big{|}D(\hat{\mu}_{n},\hat{\nu}_{n})-D(\mu,\nu)\big{|}\leq K\quad\text{almost surely}$
for a suitable constant $K$.
The proof for the compactly supported case follows from applying McDiarmid’s inequality [McD89]. When the supports are unbounded, however, the bounded difference condition does not hold, and McDiarmid’s inequality is not directly applicable. To overcome this, we will use the following version of McDiarmid’s inequality, due essentially to [Com24], tailored to our use case, which only requires the bounded difference condition to hold on a high-probability event.
Lemma 3.1 (A version of McDiarmid’s inequality).
Let $Z_{1},\dots,Z_{N}$ be independent random variables with each $Z_{i}$ taking values in a measurable space $\mathcal{Z}_{i}$. Let $\mathfrak{f}:\prod_{i=1}^{N}\mathcal{Z}_{i}\to\mathbb{R}_{+}$ be a nonnegative measurable function, for which there exist a measurable subset $\mathcal{W}\subset\prod_{i=1}^{N}\mathcal{Z}_{i}$ and nonnegative constants $\mathsf{c}_{1},\dots,\mathsf{c}_{N}$ such that
$|\mathfrak{f}(z)-\mathfrak{f}(z^{\prime})|\leq\sum_{i=1}^{N}\mathsf{c}_{i}\mathbbm{1}_{\{z_{i}\neq z_{i}^{\prime}\}},\quad z,z^{\prime}\in\mathcal{W}.$
Let $Z=(Z_{1},\dots,Z_{N})$. Assume $\mathbb{E}[\mathfrak{f}(Z)]$ is finite and $\mathfrak{p}:=\mathbb{P}(Z\notin\mathcal{W})\leq\frac{1}{2}$. Then, we have
$\mathbb{P}\left(\mathfrak{f}(Z)\geq 2\mathbb{E}[\mathfrak{f}(Z)]+t\sqrt{\sum_{i-1}^{N}\mathsf{c}_{i}^{2}/2}+\mathfrak{p}\sum_{i=1}^{N}\mathfrak{c}_{i}\right)\leq\mathfrak{p}+e^{-t^{2}},\quad t>0.$
The original formulation of Proposition 2 in [Com24] involves the conditional expectation $\mathbb{E}[\mathfrak{f}(Z)\mid Z\in\mathcal{W}]$ in place of $2\mathbb{E}[\mathfrak{f}(Z)]$. Nonnegativity of $\mathfrak{f}$ and the assumption that $\mathfrak{p}\leq\frac{1}{2}$ allows to replace the conditional mean $\mathbb{E}[\mathfrak{f}(Z)\mid Z\in\mathcal{W}]$ with $2\mathbb{E}[\mathfrak{f}(Z)]$, which is more convenient for our purpose.
@@ 4. Proofs
We will use the following observation without further mentioning: $\mathfrak{m}_{p}(\mu)\leq 1+\mathfrak{m}_{p^{\prime}}(\mu)$ for any $p<p^{\prime}$.
@@
$\sup_{\begin{subarray}{c}(\mu,\nu):~\mathfrak{m}_{4q}(\mu)\vee\mathfrak{m}_{4q}(\nu)\leq M\\ \|A\|_{\operatorname{op}}\leq R\end{subarray}}\mathbb{E}\left[\big{|}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{|}\right]\lesssim\varphi_{n,m},$
where the hidden constant depends only on $d_{x},d_{y},q,M$ and $R$.
@@
The proof is essentially contained in the proof of Theorem 4.2 in [ZGMS24], so we only provide an outline. In this proof, the notation $\lesssim$ means that an inequality holds up to a constant that depends only on $d_{x},d_{y}$ and $R$.
For any random vector $(X,Y)$ supported in $B_{\mathcal{X}}(0,r)\times B_{\mathcal{Y}}(0,r)$,
@@
From $\|\cdot\|_{\infty}$-entropy number estimates for $\mathcal{F}$ (cf. Corollary 2.7.10 in [vdVW96]), combined with a version of Dudley’s entropy integral bound (cf. Theorem 16 in [vLB04]), we have
$\mathbb{E}[I]\lesssim\inf_{\alpha>0}\left\{\alpha+\frac{1}{\sqrt{n}}\int_{\alpha}^{1}\tau^{-d_{x}/4}\,d\tau\right\}\lesssim n^{-\frac{2}{d_{x}\vee 4}}(\log n)^{\mathbbm{1}_{\{d_{x}=4\}}}.$
For the second term $II$, by Lemma 2.1 in [HSM24] (or by the definition of the $c_{A}$-transform), $\|\cdot\|_{\infty}$-covering numbers for $\mathcal{F}^{c_{A}}$ are no greater than those for $\mathcal{F}$, and as such, arguing as in the previous case, we have $\mathbb{E}[II]\lesssim m^{-\frac{2}{d_{x}\vee 4}}(\log m)^{\mathbbm{1}_{\{d_{x}=4\}}}$. The conclusion follows from the observation that
$n^{-\frac{2}{d_{x}\vee 4}}(\log n)^{\mathbbm{1}_{\{d_{x}=4\}}}+m^{-\frac{2}{d_{x}\vee 4}}(\log m)^{\mathbbm{1}_{\{d_{x}=4\}}}\lesssim\varphi_{n,m}.$
Indeed, this is trivial except for $d_{x}=4$, so assume $d_{x}=4$ and that $n\leq m$ without loss of generality. Setting $C=m/n\geq 1$, we have $m^{-1/2}\log m=n^{-1/2}C^{-1/2}(\log n+\log C)$, and since the function $z\mapsto z^{-1/2}\log z$ is bounded on $1,\infty)$, we have $m^{-1/2}\log m\lesssim n^{-1/2}\log n$. ∎
@@
$\begin{split}&\mathbb{E}\left[\big{|}T_{\bar{c}_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{\bar{c}_{A}}(\mu,\nu)-T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})+T_{c_{A}}(\mu,\nu)\big{|}\right]\\ &\leq\mathbb{E}\Big{[}2\big{|}\mathfrak{m}_{4}(\hat{\mu}_{n})-\mathfrak{m}_{4}(\mu)\big{|}+2\big{|}\mathfrak{m}_{4}(\hat{\nu}_{m})-\mathfrak{m}_{4}(\nu)\big{|}\\ &\qquad+16R\big{|}\mathfrak{m}_{2}(\hat{\mu}_{n})-\mathfrak{m}_{2}(\mu)\big{|}+16R\big{|}\mathfrak{m}_{2}(\hat{\nu}_{m})-\mathfrak{m}_{2}(\nu)\big{|}\Big{]}\\ &\leq\left\{2\sqrt{\mathfrak{m}_{8}(\mu)}+2\sqrt{\mathfrak{m}_{8}(\nu)}+16R\sqrt{\mathfrak{m}_{4}(\mu)}+16R\sqrt{\mathfrak{m}_{4}(\nu)}\right\}(n\wedge m)^{-1/2}.\end{split}$
@@
$\begin{split}\bar{c}_{A}(x,y)&\leq(4\|x\|^{4}+32R\|x\|^{2})+(4\|y\|^{4}+32R\|y\|^{2})\\ &\leq\underbrace{\big{\{}(4+32R)\|x\|^{4}+32R\big{\}}}_{=:\bar{c}_{\mathcal{X}}(x)}+\underbrace{\big{\{}(4+32R)\|y\|^{4}+32R\big{\}}}_{=:\bar{c}_{\mathcal{Y}}(y)},\end{split}$
and $\|\bar{c}_{\mathcal{X}}\|_{L^{q}(\mu)}^{q}\vee\|\bar{c}_{\mathcal{Y}}\|_{L^{q}(\nu)}^{q}\lesssim 1$ up to a constant that depends only on $q,M$ and $R$. Since
$\bar{c}_{\mathcal{X}}^{-1}([0,r])=\begin{cases}B_{\mathcal{X}}\left(0,\left(\frac{r-32R}{4+32R}\right)^{1/4}\right)&r\geq 32R,\\ \varnothing&r<32R,\end{cases}$
Condition (30) holds with $\alpha=\frac{2}{(d_{x}\wedge d_{y})\vee 4}$ and $\delta=\mathbbm{1}_{\{d_{x}\wedge d_{y}=4\}}$. The desired result then follows from Theorem A.1. ∎
@@
Let $q,R\geq 1$ be given. Set $\bar{c}_{\mathcal{X}}(x):=(4+32R)\|x\|^{4}+32R$ and $\bar{c}_{\mathcal{Y}}(y):=(4+32R)\|y\|^{4}+32R$. For every $(\mu,\nu)\in\mathcal{P}_{4q}(\mathcal{X})\times\mathcal{P}_{4q}(\mathcal{Y})$ and $A\in\mathbb{R}^{d_{x}\times d_{y}}$ with $\|A\|_{\operatorname{op}}\leq R$, one can find dual potentials $(f_{A},g_{A})$ for $(\mu,\nu)$ w.r.t. cost $c_{A}$ such that
@@
$f_{A}(x):=\bar{f}_{A}(x)-(2\|x\|^{4}+16R\|x\|^{2})\quad\text{and}\quad g_{A}(y):=\bar{g}_{A}(y)-(2\|y\|^{4}+16R\|y\|^{2})$
@@
Pick any $(\mu,\nu)\in\mathcal{P}_{4q}(\mathcal{X})\times\mathcal{P}_{4q}(\mathcal{Y})$ with $\mathfrak{m}_{4q}(\mu)\vee\mathfrak{m}_{4q}(\nu)\leq M$. In this proof, the notation $\lesssim$ means that an inequality holds up to a constant that depends only on $d_{x},d_{y},q$ and $M$. Assume without loss of generality that $\mu$ and $\nu$ have mean zero. Let $\tilde{\mu}_{n}$ and $\tilde{\nu}_{m}$ be the centered versions of $\hat{\mu}_{n},\hat{\nu}_{m}$, respectively. We first observe that
@@
These estimates follow from tedious but straightforward computations; for completeness, we provide their proofs in Lemma 4.4 below.
Let $A^{\star}$ be a matrix with $\|A^{\star}\|_{\operatorname{op}}\leq\sqrt{\mathfrak{m}_{2}(\mu)\mathfrak{m}_{2}(\nu)}/2$ that achieves the infimum in the variational representation (6),
@@
$\mathcal{E}_{n,m}=\big{\{}\mathfrak{m}_{2}(\hat{\mu}_{n})\leq\mathfrak{m}_{2}(\mu)+1\big{\}}\cap\{\mathfrak{m}_{2}(\hat{\nu}_{m})\leq\mathfrak{m}_{2}(\nu)+1\big{\}}.$
By Chebyshev’s inequality, we have
$\begin{split}\mathbb{P}(\mathcal{E}_{n,m}^{c})&\leq\mathbb{P}\big{(}\mathfrak{m}_{2}(\hat{\mu}_{n})-\mathfrak{m}_{2}(\mu)\geq 1\big{)}+\mathbb{P}\big{(}\mathfrak{m}_{2}(\hat{\nu}_{m})-\mathfrak{m}_{2}(\nu)\geq 1\big{)}\\ &\leq\big{(}\mathfrak{m}_{4}(\mu)+\mathfrak{m}_{4}(\nu)\big{)}(n\wedge m)^{-1}.\end{split}$
Observe that on the event $\mathcal{E}_{n,m}$, by Lemma 2.2, the following variational representation holds:
@@
where $\mathcal{A}:=\big{\{}A\in\mathbb{R}^{d_{x}\times d_{y}}:\|A\|_{\operatorname{op}}\leq R\big{\}}$ with $R:=\sqrt{(\mathfrak{m}_{2}(\mu)+1)(\mathfrak{m}_{2}(\nu)+1)}/2$. An analogous variational representation holds for $S_{2}(\mu,\nu)$. As such, on the event $\mathcal{E}_{n,m}$,
@@
$\begin{split}|c_{A}(x,y)-c_{B}(x,y)|&\leq 32\|x\|\|y\|\|A-B\|_{\operatorname{op}}\\ &\leq 16(\|x\|^{2}+\|y\|^{2})\|A-B\|_{\operatorname{op}}.\end{split}$
For $\epsilon>0$, let $\mathcal{N}_{\epsilon}$ be an $\epsilon$-net for $\mathcal{A}$ w.r.t. $\|\cdot\|_{\operatorname{op}}$, so that for any $A\in\mathcal{A}$, there exists $B\in\mathcal{N}_{\epsilon}$ such that $\|A-B\|_{\operatorname{op}}\leq\epsilon$. By volumetric argument, it is not difficult to see that
@@
$\begin{split}\inf_{A\in\mathcal{A}}\big{\{}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{\}}&\geq\min_{A\in\mathcal{N}_{\epsilon_{n,m}}}\big{\{}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{\}}\\ &\quad-16\big{(}\mathfrak{m}_{2}(\hat{\mu}_{n})+\mathfrak{m}_{2}(\hat{\nu}_{m})+\mathfrak{m}_{2}(\mu)+\mathfrak{m}_{2}(\nu)\big{)}\epsilon_{n,m}.\end{split}$
@@
Outside $\mathcal{E}_{n,m}$, we use the following crude bound:
$S_{2}(\hat{\mu}_{n},\hat{\nu}_{m})-S_{2}(\mu,\nu)\geq S_{2}(\hat{\mu}_{n},\hat{\nu}_{m})\geq-2\mathfrak{m}_{4}(\hat{\mu}_{n})-4\mathfrak{m}_{2}^{2}(\hat{\mu}_{n})-2\mathfrak{m}_{4}(\hat{\nu}_{m})-4\mathfrak{m}_{2}^{2}(\hat{\nu}_{m}).$
@@
$\begin{split}\mathbb{E}\left[\big{|}S_{2}(\hat{\mu}_{n},\hat{\nu}_{m})-S_{2}(\mu,\nu)\big{|}\right]&\lesssim\mathbb{E}\left[\big{|}T_{c_{A^{\star}}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A^{\star}}}(\mu,\nu)\big{|}\right]\\ &\qquad+\mathbb{E}\left[\max_{A\in\mathcal{N}_{\epsilon_{n,m}}}\left|\int f_{A}\,d(\hat{\mu}_{n}-\mu)\right|\right]\\ &\qquad+\mathbb{E}\left[\max_{A\in\mathcal{N}_{\epsilon_{n,m}}}\left|\int g_{A}\,d(\hat{\nu}_{m}-\nu)\right|\right]\\ &\qquad+\mathbb{E}\big{[}\big{(}\mathfrak{m}_{4}(\hat{\mu}_{n})+\mathfrak{m}_{4}(\hat{\nu}_{m})\big{)}\mathbbm{1}_{\mathcal{E}_{n,m}^{c}}\big{]}+(n\wedge m)^{-\frac{2}{(d_{x}\wedge d_{y})\vee 4}}.\end{split}$